Shift summary for evening shift 8/13/99

Robert Tschirhart (tsch@fnal.gov)
Sat, 14 Aug 1999 00:00:20 -0500

Shift Summary for owl shift, 07/16/99

On Shift: R. Barbosa, A. Soddu, BobT

Hours of beam available at all intensities: 4 hours
Hours of beam available at nominal intensity: 3 hours
Hours of E832 data (special runs not included): 4 hours
Hours of special runs (laser scans, muon runs...): 0 hours

Accounting of lost time: Spotty beam, Booster vacuum/TeV problems.

========================================================================

Runs ended this shift:

Run 14223 - BEAM_832 (2.6 M events)

Tapes brought to FCC this shift: none
========================================================================

Accesses during shift: none
========================================================================

Overview of shift: Inherited fairly smooth 832 running at 6.8E12 from the
day shift, which became spotty at the begining of the
evening shift. Beam downtime was used to verify CsI
HV settings with respect to the database.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14223 17:10 No Beam, none expected for a while. We end
this run and verify the measured CsI HV with
respect to the database. All channels verify
correctly except CsI 1136, where the measured
HV is correct, but the database is low by 116
volts.

14224 17:50 Beam returns.


14224 51 18:35 Autotune locks up, and beam is found wandering
6-7 mm off vertical nominal in SEED2. First
50 spills of 14224 suffer some beam wandering.
We contact MCR, and they fix it.

14224 19:30 Targeting ok, but now SEM is wandering high.
Lots of spills at 7.0E12-7.3E12. Discussions
with MCR bring it back down.

14224 20:00
- 22:00 Spotty beam. Discovered a bad point in the
E/P distribution of CsI 1773. No shift in the
laser gain. This is very likely just one
bad point, need to keep an eye on it.

14224 22:00 No beam.
- 24:00

========================================================================
Follow up Notes on Previous Shift

Verify CsI HV wrt database, done.

========================================================================
Items to Follow up on Next Shift

Keep an eye on CsI 1773 in csigain#list, probably suffering from a
single bad data point.